# Reassessing quantum-thermodynamic enhancements in continuous thermal machines

### v2

@@ II.1 Coherent three-level amplifier
The three-level maser or amplifier, as initially introduced by Scovil and Schulz-DuBois in Ref. [3], is one of the simplest models of a thermal machine, capable of serving either as a heat engine or a refrigerator depending on the configuration of system parameters [45, 46]. The characteristics and performance of this model has been largely studied [4, 46, 5, 15, 47, 48, 1, 49, 30] owing to its simplicity and versatile functionality, and a first experimental implementation has been reported in Ref. [12]. It stands as a main example of a thermal machine where coherence among non-degenerate energy levels is supported in the steady state, suggesting the appearance of regimes where quantum-enhanced thermodynamic operation can be achieved [9] (for a critical view of such quantum-enhanced performance see e.g. Ref. [23]).
@@
The two remaining transitions of the three-level system are further weakly coupled to two thermal baths at different inverse temperatures, denoted as cold and hot ($\beta_{\mathrm{c}}\geq\beta_{\mathrm{h}}$), typically modeled as bosonic reservoirs. They lead to four incoherent jumps described by Lindblad operators:
@@
with rates $\gamma_{21}=\gamma_{\mathrm{c}}(\bar{n}_{\mathrm{c}}+1)$ and $\gamma_{12}=\gamma_{\mathrm{c}}\bar{n}_{\mathrm{c}}$ associated respectively to the emission and absorption of energy quanta $\Delta\epsilon_{\mathrm{c}}=\epsilon_{2}-\epsilon_{1}$ into the cold reservoir at $\beta_{\mathrm{c}}$. Similarly $\gamma_{20}=\gamma_{\mathrm{h}}(\bar{n}_{\mathrm{h}}+1)$ and $\gamma_{02}=\gamma_{\mathrm{h}}\bar{n}_{\mathrm{h}}$ stand for emission and absorption of energy quanta $\Delta\epsilon_{\mathrm{h}}=\epsilon_{2}$ into the hot reservoir at temperature $\beta_{\mathrm{h}}$. Here $\gamma_{r}$ denotes the spontaneous emission rate for bath $r=\mathrm{c,h}$ and $\bar{n}_{r}=(e^{\beta_{r}\Delta\epsilon_{r}}-1)^{-1}$ is the average number of thermal photons with frequency $\Delta\epsilon_{r}$ in the baths.
@@
II.2 Three-qubit autonomous refrigerator
This fridge model was first presented in Ref. [6] and consist of one of the smallest thermal machines models using a multipartite system (see Fig. 1b). It pertains to the class of autonomous quantum refrigerators [50, 8, 51], also called quantum absorption refrigerators [7, 52, 53, 54, 21, 55]. It consists of three qubits with different energy gaps, each of them locally coupled to a corresponding thermal baths at a different temperature. A weak three-body energy-preserving Hamiltonian interaction among the qubits allows the generation of heat currents between the three reservoirs that ultimately power thermodynamic tasks such as heat pumping or refrigeration [8]. Different platforms for the actual implementation of this or closely related models in the lab has been proposed using quantum dots [56], optical systems [57], QED arquitectures [58] and trapped ions [11]. In this multipartite setup, genuine quantum features such as quantum entanglement [20] and discord [52] have been analyzed, suggesting the possibility of boosting the fridge performance by quantum correlations in some operational regimes [20].
@@ II.2 Three-qubit autonomous refrigerator
Importantly, by assuming the resonance condition $\epsilon_{3}=\epsilon_{2}-\epsilon_{1}$, the three qubit interaction verifies strict energy preservation between the qubits, that is $[V,H_{1}+H_{2}+H_{3}]=0$, ensuring that the energy exchanges among the fridge qubits occur without the need of any extra source of energy or control, i.e. preserving the autonomy of the model. At difference from other autonomous fridges, such as single-qutrit fridges [46], this model exhibits steady-state coherence between degenerate energy levels |$\ket{101}$ and |$\ket{010}$ due to the presence of the interaction $V$ [6], which ultimately leads to entanglement among different partitions involving the qubits [20].
In this case all the transitions are mediated by the reservoirs, with either cold, medium or hot temperatures ($\beta_{\mathrm{c}}\geq\beta_{\mathrm{m}}\geq\beta_{\mathrm{h}}$). Since each qubit $i$ is locally coupled only to a single bath at inverse temperature $\beta_{i}$, and the interaction $V$ is weak, the master equation (2) adopts a local form [44, 43, 42], with six incoherent jumps described by Lindblad operators promoting local jumps in each qubit:
@@
together with the opposite jumps, $L_{\uparrow}^{(r)}=e^{-\beta_{r}\Delta\epsilon_{r}/2}L_{\downarrow}^{(r)\dagger}$, for $r=\text{c},\text{m},\text{h}$. The rates $\gamma_{10}^{\text{c}}=\gamma_{\mathrm{c}}(\bar{n}_{\mathrm{c}}+1)$ and $\gamma_{01}^{\text{c}}=\gamma_{\mathrm{c}}\bar{n}_{\mathrm{c}}$ are associated respectively to the emission and absorption of energy quanta $\Delta\epsilon_{\mathrm{c}}=\epsilon_{1}$ into the cold reservoir at $\beta_{\mathrm{c}}$. Similarly we have $\gamma_{10}^{\text{m}}=\gamma_{\mathrm{m}}(\bar{n}_{\mathrm{m}}+1)$ and $\gamma_{01}=\gamma_{\mathrm{m}}\bar{n}_{\mathrm{m}}$ for the emission and absorption of energy quanta $\Delta\epsilon_{\mathrm{m}}=\epsilon_{2}$ into the medium reservoir at temperature $\beta_{\mathrm{m}}$, as well as $\gamma_{10}^{\text{h}}=\gamma_{\mathrm{h}}(\bar{n}_{\mathrm{h}}+1)$ and $\gamma_{01}^{\text{h}}=\gamma_{\mathrm{h}}\bar{n}_{\mathrm{h}}$ for the emission and absorption of energy quanta $\Delta\epsilon_{\mathrm{h}}=\epsilon_{3}$ into the hot reservoir at temperature $\beta_{\mathrm{h}}$.
II.3 Noise-induced-coherence machine
A final set of continuous thermal machines models showing quantum effects, which we collectively dub noise-induced-coherence (NIC) machines, were first presented in a series of papers by Scully et al. [59, 16, 19]. In these machines, degenerate levels in the energy spectrum are combined with a collective action of the baths on the system transitions to generate coherence in the steady state [60, 61]. The operation and performance of these kind of machines has been extensively investigated within the context of quantum thermal machines [62, 63, 17, 64, 65, 48, 66, 22, 67, 68, 18], pointing to enhancements in the power output for adequate operational regimes.
A notable example of this coherence is found in a 4-level absorption refrigerator than can also work as a heat engine, whose thermodynamics have been examined in previous studies [66, 69, 68]. In this system, two of the levels possess the same energy, as depicted in Fig. 1c. These levels are subjected to the influence of two distinct thermal baths, whose action results in the emergence of degenerate coherence that persist in the system’s steady state. The system’s Hamiltonian in this case is as follows
@@ II.3 Noise-induced-coherence machine
with rates $\gamma_{i1}=\gamma_{\mathrm{c}}^{i}(\bar{n}_{\mathrm{c}}+1)$ and $\gamma_{1i}=\gamma_{\mathrm{c}}^{i}\bar{n}_{\mathrm{c}}$ for $i=\text{a},\text{b}$, associated respectively to the emission and absorption of energy quanta $\Delta\epsilon_{\mathrm{c}}=\epsilon_{2}-\epsilon_{1}$ into the cold reservoir at $\beta_{\mathrm{c}}$, and similarly $\gamma_{i0}=\gamma_{\mathrm{h}}^{i}(\bar{n}_{\mathrm{h}}+1)$ and $\gamma_{0i}=\gamma_{\mathrm{h}}^{i}\bar{n}_{\mathrm{h}}$ for $i=\text{a},\text{b}$, is associated with emission and absorption of energy quanta $\Delta\epsilon_{\mathrm{h}}=\epsilon_{2}$ into the hot reservoir at temperature $\beta_{\mathrm{h}}$. Finally, since the “work” bath is at an infinite temperature i.e $\beta_{\text{w}}\rightarrow 0$ the rates associated to it satisfy $\gamma_{10}=\gamma_{01}$, corresponding in this case for emission and absorption of a quanta $\Delta\epsilon_{\text{w}}=\epsilon_{1}$ from the work source.
III Thermodynamic performance
We are interested in the performance of the quantum thermal machine models presented in the previous section when operating in non-equilibrium steady-state conditions [15, 1]. By performance, we refer not only to the size of the output current generated by the machine operation (output power in the case of heat engines or cooling power for the case of refrigerators) and the machine thermodynamic efficiency (ratio of useful output to source input), but also to the size of the fluctuations in the output current, which can be viewed as measure of the “quality” of that output in stochastic machines [27].
@@ III Thermodynamic performance
$\langle\dot{Q}_{r}\rangle$ $:=\sum_{k}\operatorname{Tr}[H_{0}\mathcal{D}_{k}^{(r)}[\pi(t)]],$ (13)
where we recall that $\pi(t)$ may acquire a periodic time-dependence (in Schrödinger picture) due to the presence of non-diagonal elements (coherences) in the steady-state density operator. We also emphasize that for weak perturbations as the ones considered here, only the bare Hamiltonian $H_{0}$ enters in the heat currents [71, 72], ensuring consistency with the second law [73] (see also Appendix A). The first law takes the form $\langle\dot{W}\rangle=\sum_{r}\langle\dot{Q}_{r}\rangle$, imposing that any output power is sustained by input heat currents from the baths. Explicit expressions of the heat currents $\langle\dot{Q}_{r}\rangle$ valid for generic machines (with or without degeneracy) are given in Appendix A.
@@
which characterizes the irreversibility of the machine operation in its nonequilibrium steady-state [71, 2]. Here it is also work remarking that whenever the temperature of some of the reservoirs $r$ approaches infinity, $\beta_{r}\rightarrow 0$, the associated energy current does not contribute to the entropy production and hence it should be considered as (incoherent) work rather than heat, see also Refs. [7, 74]. In that case the output power associated to such a work reservoir reads $\langle\dot{W}\rangle=-\sum_{k}\operatorname{Tr}[H_{0}\mathcal{D}_{k}^{(r)}[\pi(t)]]$.
@@
where in the case of heat engine operation ${J}_{\mathrm{out}}=\langle\dot{W}\rangle$ and ${J}_{\mathrm{in}}=\langle\dot{Q}_{\mathrm{h}}\rangle$, while for refrigeration the efficiency (coefficient of performance) is given from ${J}_{\mathrm{out}}=\langle\dot{Q}_{\mathrm{c}}\rangle$ and either ${J}_{\mathrm{in}}=-\langle\dot{W}\rangle$ for power-driven refrigerators (as the models in Secs. II.1 and II.3), or ${J}_{\mathrm{in}}=-\langle\dot{Q}_{\mathrm{h}}\rangle$ for absorption refrigerators (as the one in Sec. II.2). For extensions of efficiency to multiple inputs and outputs see e.g. Refs. [75, 76, 77].
By combining the first and second laws in the setup we recover Carnot bound for the efficiency of heat engines $\eta\leq\eta_{\mathrm{C}}:=1-{\beta_{\mathrm{h}}/\beta_{\mathrm{c}}}$, as well as the corresponding (Carnot) bounds for power-driven fridges $\eta\leq\eta_{\mathrm{fr}}:=\beta_{\mathrm{h}}/(\beta_{\mathrm{c}}-\beta_{\mathrm{h}})$ and absorption refrigerators, $\eta\leq\eta_{\mathrm{abs}}:=(\beta_{\mathrm{m}}-\beta_{\mathrm{h}})/(\beta_{\mathrm{c}}-\beta_{\mathrm{m}})$, achieved in the limit of zero entropy production, where all energy currents vanish [15, 1, 50].
The maximum efficiency (zero power) equilibrium points separate the different modes of operation of the machines, where heat currents and output power change sign. In the models presented here they can be easily obtained analytically by noticing that the average steady-state currents can be rewritten as:
@@
where $\langle\dot{N}_{r}\rangle$ is the probability current associated to the reservoir $r$ (flux of quanta). It results from the fact that in the models, baths are associated to a single energy gap $\Delta\epsilon_{r}$. In the steady state, due to the absence of leaks, the system will exchange excitations with all baths at the same rate, i.e. $\forall i,j$ we have $\langle\dot{N}_{i}\rangle=\langle\dot{N}_{j}\rangle:=\langle\dot{N}\rangle$. In addition, the first law constrains the average value of the work to satisfy $\langle\dot{W}\rangle=\sum_{r}\Delta\epsilon_{r}\langle\dot{N}\rangle=\omega_{d}\langle\dot{N}\rangle$. The ratio between the different steady-state currents hence verify
$\frac{|\langle\dot{W}\rangle|}{|\langle\dot{Q}_{j}\rangle|}=\frac{\omega_{\mathrm{d}}}{\epsilon_{j}}~{}~{}~{};~{}~{}~{}\frac{|\langle\dot{Q}_{i}\rangle|}{|\langle\dot{Q}_{j}\rangle|}=\frac{\epsilon_{i}}{\epsilon_{j}}.$ (17)
By combining the above proportionality relations with the efficiency bounds above, we can construct a table for the operational modes of each model (see Table 1).
@@
While it is in general desirable for any thermal machine to have a large output current and a high efficiency (low rate of entropy production), in microscopic systems another fundamental factor to consider is the fluctuations associated with the currents, specially in the output power. Therefore, in addition to minimizing entropy production and maximizing power, low fluctuations (resulting in higher precision) in energy flows are also desirable. In classical systems these three quantities are not independent, but their trade-off is quantified by the TUR:
@@
where in this case we identified $J_{\mathrm{out}}=\dot{Q}_{\mathrm{c}}$ and the entropy production rate reads $\langle\dot{S}_{\mathrm{tot}}\rangle=-\beta_{\mathrm{h}}\langle\dot{Q}_{\mathrm{h}}\rangle-\beta_{\mathrm{c}}\langle\dot{Q}_{\mathrm{c}}\rangle-\beta_{\mathrm{m}}\langle\dot{Q}_{\mathrm{m}}\rangle=(\beta_{\mathrm{c}}-\beta_{\mathrm{m}})(\eta_{\mathrm{abs}}/\eta-1)\langle\dot{Q}_{\mathrm{c}}\rangle$. Throughout this paper, to evaluate fluctuations in the output currents, i.e. the variances $\text{Var}[\dot{W}]$ and $\text{Var}[\dot{Q}_{\mathrm{c}}]$, we employ the full-counting statistics (FCS) formalism [78, 79] (see details in Appendix B).

### v4

@@ A.1 Coherent three-level amplifier
The three-level maser or amplifier, as initially introduced by Scovil and Schulz-DuBois in Ref. [3], is one of the simplest models of a thermal machine, capable of performing either as a heat engine or a refrigerator, depending on the configuration of system parameters [66, 67]. The characteristics and performance of this model have been largely studied [4, 67, 5, 15, 46, 68, 1, 69, 30] owing to its simplicity and versatile functionality, and a first experimental implementation has been reported in Ref. [12]. It stands as a main example of a thermal machine where coherence among non degenerate energy levels is supported in the steady state, suggesting the appearance of regimes where quantum-enhanced thermodynamic operation can be achieved [9] (for a critical view of such quantum-enhanced performance, see e.g. Ref. [23]).
@@
The two remaining transitions of the three-level system are weakly coupled to two thermal baths at different inverse temperatures, denoted as cold and hot ($\beta_{\mathrm{c}}\geq\beta_{\mathrm{h}}$), typically modeled as bosonic reservoirs. They lead to four incoherent jumps described by Lindblad operators:
@@
with rates $\gamma_{21}=\gamma_{\mathrm{c}}(\bar{n}_{\mathrm{c}}+1)$ and $\gamma_{12}=\gamma_{\mathrm{c}}\bar{n}_{\mathrm{c}}$ associated respectively to the emission and absorption of energy quanta $\Delta\epsilon_{\mathrm{c}}=\epsilon_{2}-\epsilon_{1}$ into the cold reservoir at $\beta_{\mathrm{c}}$. Similarly $\gamma_{20}=\gamma_{\mathrm{h}}(\bar{n}_{\mathrm{h}}+1)$ and $\gamma_{02}=\gamma_{\mathrm{h}}\bar{n}_{\mathrm{h}}$ stand for emission and absorption of energy quanta $\Delta\epsilon_{\mathrm{h}}=\epsilon_{2}$ into the hot reservoir at temperature $\beta_{\mathrm{h}}$. Here, $\gamma_{r}$ denotes the spontaneous emission rate for bath $r=\mathrm{c,h}$ and $\bar{n}_{r}=(e^{\beta_{r}\Delta\epsilon_{r}}-1)^{-1}$ is the average number of thermal photons with frequency $\Delta\epsilon_{r}$ in the baths.
@@
A.2 Three-qubit autonomous refrigerator
This fridge model was first presented in Ref. [6] and consists of one of the smallest thermal machine models using a multipartite system (see Fig. 1b). It pertains to the class of autonomous quantum refrigerators [70, 8, 71], also called quantum absorption refrigerators [7, 72, 64, 73, 21, 74]. It consists of three qubits with different energy gaps, each of them locally coupled to a corresponding thermal bath at a different temperature. A weak three-body energy-preserving Hamiltonian interaction among the qubits allows the generation of heat currents between the three reservoirs that ultimately power thermodynamic tasks such as heat pumping or refrigeration [8]. Different platforms for the actual implementation of this or closely related models in the lab, have been proposed using quantum dots [44], optical systems [54], QED architectures [53] and trapped ions [11]. In this multipartite setup, genuine quantum features such as quantum entanglement [20] and discord [72] have been analyzed, suggesting the possibility of boosting the fridge performance by quantum correlations in some operational regimes [20].
@@ A.2 Three-qubit autonomous refrigerator
Importantly, by assuming the resonance condition $\epsilon_{3}=\epsilon_{2}-\epsilon_{1}$, the three-qubit interaction verifies strict energy preservation between the qubits, that is, $[V,H_{1}+H_{2}+H_{3}]=0$, ensuring that the energy exchanges among the fridge qubits occur without the need of any extra source of energy or control, i.e. preserving the autonomy of the model. At difference from other autonomous fridges, such as single-qutrit fridges [67], this model exhibits steady-state coherence between degenerate energy levels |$\ket{101}$ and |$\ket{010}$ due to the presence of interaction $V$ [6], which ultimately leads to entanglement among different partitions involving the qubits [20].
In this case, all transitions are mediated by the reservoirs, with either cold, mild, or hot temperatures ($\beta_{\mathrm{c}}\geq\beta_{\mathrm{m}}\geq\beta_{\mathrm{h}}$). Since each qubit $i$ is locally coupled only to a single bath at inverse temperature $\beta_{i}$, and the interaction $V$ is weak, the master equation (2) adopts a local form [38, 37, 36], with six incoherent jumps described by Lindblad operators promoting local jumps in each qubit:
@@
together with the opposite jumps, $L_{\uparrow}^{(r)}=e^{-\beta_{r}\Delta\epsilon_{r}/2}L_{\downarrow}^{(r)\dagger}$, for $r=\text{c},\text{m},\text{h}$. The rates $\gamma_{10}^{\text{c}}=\gamma_{\mathrm{c}}(\bar{n}_{\mathrm{c}}+1)$ and $\gamma_{01}^{\text{c}}=\gamma_{\mathrm{c}}\bar{n}_{\mathrm{c}}$ are associated, respectively, to the emission and absorption of energy quanta $\Delta\epsilon_{\mathrm{c}}=\epsilon_{1}$ into the cold reservoir at $\beta_{\mathrm{c}}$. Similarly, we have $\gamma_{10}^{\text{m}}=\gamma_{\mathrm{m}}(\bar{n}_{\mathrm{m}}+1)$ and $\gamma_{01}=\gamma_{\mathrm{m}}\bar{n}_{\mathrm{m}}$, for the emission and absorption of energy quanta $\Delta\epsilon_{\mathrm{m}}=\epsilon_{2}$ into the medium reservoir at temperature $\beta_{\mathrm{m}}$, as well as $\gamma_{10}^{\text{h}}=\gamma_{\mathrm{h}}(\bar{n}_{\mathrm{h}}+1)$ and $\gamma_{01}^{\text{h}}=\gamma_{\mathrm{h}}\bar{n}_{\mathrm{h}}$ for the emission and absorption of energy quanta $\Delta\epsilon_{\mathrm{h}}=\epsilon_{3}$ into the hot reservoir at temperature $\beta_{\mathrm{h}}$.
A.3 Noise-induced-coherence machine
A final set of continuous thermal machine models showing quantum effects, which we collectively dub noise-induced-coherence (NIC) machines, were first presented in a series of papers by Scully et al. [75, 16, 19]. In these machines, degenerate levels in the energy spectrum are combined with a collective action of the baths on the system transitions to generate coherence in the steady state [76, 77]. The operation and performance of these kind of machines has been extensively investigated within the context of quantum thermal machines [78, 79, 17, 80, 81, 68, 58, 22, 82, 57, 18], pointing to output power enhancements for adequate operational regimes.
A notable example of this coherence is found in a 4-level absorption refrigerator that can also work as a heat engine, whose thermodynamics has been examined in previous studies [58, 60, 57]. In this system, two of the levels possess the same energy, as depicted in Fig. 1c. These levels are subjected to the influence of two distinct thermal baths whose action results in the emergence of degenerate coherence that persist in the steady state. The system’s Hamiltonian in this case is as follows
@@ A.3 Noise-induced-coherence machine
with rates $\gamma_{i1}=\gamma_{\mathrm{h}}^{i}(\bar{n}_{\mathrm{h}}+1)$ and $\gamma_{1i}=\gamma_{\mathrm{h}}^{i}\bar{n}_{\mathrm{h}}$ for $i=\text{a},\text{b}$, associated respectively to the emission and absorption of energy quanta $\Delta\epsilon_{\mathrm{h}}=\epsilon_{2}-\epsilon_{1}$ into the hot reservoir at $\beta_{\mathrm{h}}$, and similarly $\gamma_{i0}=\gamma_{\mathrm{c}}^{i}(\bar{n}_{\mathrm{c}}+1)$ and $\gamma_{0i}=\gamma_{\mathrm{c}}^{i}\bar{n}_{\mathrm{c}}$ for $i=\text{a},\text{b}$, is associated with emission and absorption of energy quanta $\Delta\epsilon_{\mathrm{c}}=\epsilon_{2}$ into the cold reservoir at temperature $\beta_{\mathrm{c}}$. Finally, since the “work” bath is at an infinite temperature i.e $\beta_{\text{w}}\rightarrow 0$ the rates associated to it satisfy $\gamma_{10}=\gamma_{01}$, corresponding in this case for emission and absorption of a quanta $\Delta\epsilon_{\text{w}}=\epsilon_{1}$ from the work source.
Appendix B Thermodynamic performance
We are interested in the performance of the quantum thermal machine models presented in the previous section, when operating in nonequilibrium steady-state conditions [15, 1]. By performance, we refer not only to the size of the output current generated by the machine operation (output power in the case of heat engines or cooling power for the case of refrigerators) and the thermodynamic efficiency of the machine (ratio of useful output to source input), but also to the size of the fluctuations in the output current, which can be viewed as a measure of the “quality” of that output in stochastic machines [27].
@@ Appendix B Thermodynamic performance
$\langle\dot{Q}_{r}\rangle$ $:=\sum_{k}\operatorname{Tr}[H\mathcal{D}_{k}^{(r)}[\pi(t)]]\simeq\sum_{k}\operatorname{Tr}[H_{0}\mathcal{D}_{k}^{(r)}[\pi(t)]],$
where we recall that $\pi(t)$ may acquire a periodic time-dependence (in Schrödinger picture) due to the presence of non-diagonal elements (coherences) in the steady-state density operator. We also emphasize that for weak perturbations $V$ as the ones considered here, in order to be consistent with the microscopic derivation of the master equation [38], only the bare Hamiltonian $H_{0}$ should enter in the heat currents, which ensures consistency with the second law [84, 59, 85] (see also App. C). The first law takes the form $\langle\dot{W}\rangle=\sum_{r}\langle\dot{Q}_{r}\rangle$, imposing that any output power is sustained by input heat currents from the baths. Explicit expressions of the heat currents $\langle\dot{Q}_{r}\rangle$ valid for generic machines (with or without degeneracy) are given in Appendix C.
@@
which characterizes the irreversibility of the machine operation in its nonequilibrium steady-state [59, 2]. Here, it is also worth remarking that, whenever the temperature of some of the reservoirs $r$ approaches infinity, $\beta_{r}\rightarrow 0$, the associated energy current does not contribute to the entropy production, and hence it should be considered as (incoherent) work rather than heat, see also Refs. [7, 86]. In that case, the output power associated to such a work reservoir reads $\langle\dot{W}\rangle=-\sum_{k}\operatorname{Tr}[H_{0}\mathcal{D}_{k}^{(r)}[\pi(t)]]$.
@@
where in the case of heat engine operation ${J}_{\mathrm{out}}=\langle\dot{W}\rangle$ and ${J}_{\mathrm{in}}=\langle\dot{Q}_{\mathrm{h}}\rangle$, while for refrigeration the efficiency (coefficient of performance) is given from ${J}_{\mathrm{out}}=\langle\dot{Q}_{\mathrm{c}}\rangle$ and either ${J}_{\mathrm{in}}=-\langle\dot{W}\rangle$ for power-driven refrigerators (as the models in Fig. 1a and 1c), or ${J}_{\mathrm{in}}=-\langle\dot{Q}_{\mathrm{h}}\rangle$ for absorption refrigerators (as the one in Fig.1b). For extensions of efficiency to multiple inputs and outputs see, e.g. Refs. [87, 88, 65].
By combining the first and second laws in the setup, we recover Carnot bound for the efficiency of heat engines $\eta\leq\eta_{\mathrm{C}}:=1-{\beta_{\mathrm{h}}/\beta_{\mathrm{c}}}$, as well as the corresponding (Carnot) bounds for power-driven fridges $\eta\leq\eta_{\mathrm{fr}}:=\beta_{\mathrm{h}}/(\beta_{\mathrm{c}}-\beta_{\mathrm{h}})$ and absorption refrigerators, $\eta\leq\eta_{\mathrm{abs}}:=(\beta_{\mathrm{m}}-\beta_{\mathrm{h}})/(\beta_{\mathrm{c}}-\beta_{\mathrm{m}})$, achieved in the limit of zero entropy production, where all energy currents vanish [15, 1, 70].
The maximum efficiency (zero power) equilibrium point separates the different modes of operation of the machines, where heat currents and output power change sign. In local dissipation models, due to the absence of heat leaks, the average steady-state currents can be rewritten as:
@@
where $\langle\dot{N}_{r}\rangle$ is the probability current associated to the reservoir $r$ (flux of quanta). This expression comes from the baths being associated with a single energy gap $\Delta\epsilon_{r}$ (the eventual action of a bath over other energy gaps may be considered as produced by an independent bath at the same temperature). In the steady state, if the system exchanges excitations with all baths at the same rate (tight coupling), we have $\forall i,j$ that $\langle\dot{N}_{i}\rangle=\langle\dot{N}_{j}\rangle:=\langle\dot{N}\rangle$. In addition, the first law constrains the average value of the work to satisfy $\langle\dot{W}\rangle=\sum_{r}\Delta\epsilon_{r}\langle\dot{N}\rangle=\omega_{d}\langle\dot{N}\rangle$. The ratio between the different steady-state currents hence verifies
$\frac{|\langle\dot{W}\rangle|}{|\langle\dot{Q}_{j}\rangle|}=\frac{\omega_{\mathrm{d}}}{\Delta\epsilon_{j}}~{}~{}~{};~{}~{}~{}\frac{|\langle\dot{Q}_{i}\rangle|}{|\langle\dot{Q}_{j}\rangle|}=\frac{\Delta\epsilon_{i}}{\Delta\epsilon_{j}},$ (37)
which is verified in the three illustrative examples. By combining the above proportionality relations with the efficiency bounds, we can construct a table for the operational modes of each model (see Table 1).
@@
In addition to minimizing entropy production and maximizing power, low fluctuations (resulting in higher precision) in energy flows are also desirable for the performance of thermal machines. In classical systems these three quantities are not independent, but their trade-off is quantified by the TUR:
@@
where in this case we identified $J_{\mathrm{out}}=\dot{Q}_{\mathrm{c}}$ and the entropy production rate reads $\langle\dot{S}_{\mathrm{tot}}\rangle=-\beta_{\mathrm{h}}\langle\dot{Q}_{\mathrm{h}}\rangle-\beta_{\mathrm{c}}\langle\dot{Q}_{\mathrm{c}}\rangle-\beta_{\mathrm{m}}\langle\dot{Q}_{\mathrm{m}}\rangle=(\beta_{\mathrm{c}}-\beta_{\mathrm{m}})(\eta_{\mathrm{abs}}/\eta-1)\langle\dot{Q}_{\mathrm{c}}\rangle$. Throughout this paper, to evaluate fluctuations in the output currents, i.e. the variances $\text{Var}[\dot{W}]$ and $\text{Var}[\dot{Q}_{\mathrm{c}}]$, we employ the full-counting statistics (FCS) formalism [89, 90].